### v1

@@ 3.2 Orientation mapping
Figure 3: (a) dark-field image of polycrystalline aluminum film. (b–e) Conventional diffraction patterns from 4 spots in (a). (f–i) Corresponding MAPED diffraction patterns show enhanced peaks at high scattering angles and reduction of Kikuchi diffraction lines. ACOM maps and dilation ($\epsilon_{xx}$ + $\epsilon_{yy}$) from (g–h) conventional and (i–j) MAPED data with (k) legend. Dark-field region in (a) marked with rectangle in (j).
To further explore the application of our MAPED method to study material structure, we performed ACOM analysis on a polycrystalline aluminum film (Fig. 3). From the dark-field image (Fig. 3a) we plot four diffraction patterns from the same position. The top row presents conventional STEM diffraction patterns, while the bottom row shows data from the same positions but collected with our MAPED method. As with the data in Fig. 2, incorporating more tilts leads to diffraction information at higher scattering angles. Moreover, the few-tilt method suppressed dynamical artifacts such as Kikuchi diffraction, which is seen comparing Fig. 3b&f.
The orientation maps and strain calculations from the ACOM maps are shown in Figs. 3j–m. The conventional diffraction data approach yields high-quality orientation matches. The in-plane rotation maps plus the out-of-plane orientation (Fig. S.8) show most of the grains within the film are aligned with the [111] parallel to the beam with in-plane rotation between grains. Finally, using the structure file as a reference, the strain in each grain is calculated, with the dilation ($\epsilon_{xx}$ + $\epsilon_{yy}$) plotted in Fig. 3k. The full strain maps are shown in Fig. S.10.
The same experiment performed on the identical field of view using MAPED yields qualitatively similar results, but with significantly improved signal-to-noise in both the ACOM maps and strain calculations (Figs. 3l–m, S.9, and S.11). The improvement is especially evident in the strain maps, where the MAPED data produces far more physical results. This trend is evident in the large grain at the bottom left of the field, highlighted by the arrow in Figs. 3j–m. While the polycrystalline aluminum film studied here serves as a model system to compare ACOM with and without precession, we anticipate that this improvement in ACOM matching will be even more significant in challenging samples.
@@ 3.3 Quantification of intensities
Figure 4: Dynamical diffraction simulations to explore impact of few tilt precession on intensity quantification. (a) Si [011] diffraction pattern. The ($\overline{2}00$) and ($\overline{11}0$) intensity values are shown in (b) and (c) as a function of thickness and number of tilts. Intensity values normalized to show percent difference from the intensity value with the highest number of tilts.
Although our MAPED method can increase the signal from high scattering angles as shown in Figs. 2&3, there is an anisotropic nature to the method, which non-uniformly boosts the signal in the diffraction patterns. For example, consider the five diffraction patterns from the silicon sample shown in Fig. 1a. The tilt of the Ewald sphere enhances the signal of some disks more than others leading to brighter reflections in the diffraction patterns.
To better quantify the impact our few-tilt precession method has on diffraction pattern intensities, we performed dynamical Bloch wave simulations in py4DSTEM. The extent of intensity error depends on sample thickness, tilt angle, and the number of tilts used. Here we simulate silicon with the [011] aligned with the electron beam, matching the experimental data in Fig. 2.
Fig. 4 shows the result of our calculations. We include one simulation at zero beam tilt, and N-1 equally spaced radially at 17 mrad. We calculated the intensities at two reflections ($\overline{2}00$) and ($\overline{11}0$), as indicated in Fig. 4a. We performed simulations for sample thicknesses from 10 to 300 nm in 10 nm steps, using 5, 9, 17, 33, 65, 129, 257, and 513 tilts. Fig. 4b shows the results for the ($\overline{2}00$) reflection. Here we normalized the intensity values to show percent difference from the intensity value with the highest number of tilts. The intensities are plotted in Fig. S.12. For the thin specimens there is little variation in intensity with number of tilts. The error in intensity increases at larger thickness values. However, the value converges as more tilts are included. One can imagine how continuously precessing systems, which more finely sample the precession space, facilitate predictable intensity values and more complete sampling of diffraction space. The $\overline{1}\overline{1}0$ reflection shows less systematic deviations. Again, thinner specimens have smaller errors in intensity and as the number of tilts increases, the intensity values converge.
@@ 4 Discussion
As discussed in the introduction, precession is well-understood to improve diffraction data by reducing dynamical diffraction artifacts and collecting higher scattering angles. Here, we have highlighted how collecting data at a few tilts can lead to high-quality diffraction data with many of the benefits of integrated precession experiments. The flexibility in choice of cameras and microscopes to collect precession data also poses a significant advantage for 4D-STEM experiments.
Our MAPED method also helps with one of the challenges in aligning precession tools for non-integrated cameras. Precession systems that are integrated with the camera address a core challenge of precession 4D-STEM: a precession-induced “descan,” which is a shift of unscattered beam in the back focal plane as a consequence of tilting the incident illumination. This precession-descan is in addition to the common descan observed when an electron beam is rastered across a specimen, and it is elegantly handled by automated routines in the integrated systems. If one does not have a commercial precession integrated with the camera of choice, it is possible to use a camera that is not integrated by having one scan generator do the X-Y scanning in synchronization with the camera and a second scan generator do precession and precession-descan correction. This mode does not have automated routines, so fine tuning of the precession-descan correction with the second scan generator is challenging. Without perfect alignment, the Bragg disks can become blurred due to precession-descan, without the ability to correct computationally after acquisition. However, in our MAPED method we are able to preserve sharp features in our diffraction disks, even in the challenging case of bullseye apertures as shown in Fig. 2, because both types of descan are addressed in post-processing.
@@
An additional consideration is the practicality of acquiring these few tilt 4D-STEM scans. Collecting and processing MAPED datasets requires additional microscope time and computational resources over conventional experiments. However, automated precession pipelines through vendor hardware also require additional overhead and ultimately require similar collection times when competing with fast cameras. An additional practical advantage of the described sequential approach is that the microscope stays in STEM mode, as compared to commercial systems, which usually use a converged probe in TEM mode. Collected data in STEM mode allows easier correlative acquisition of other microscopy signals, including annular dark-field STEM and STEM spectroscopy for multimodal analysis. Moreover, in STEM mode, we have more flexibility in convergence angles and apertures, such as the bullseye used in Fig. 2.
Despite these benefits there are limitations for the few-tilt precession approach. As studied in Sec. 3.3, especially for thicker samples, the intensities are non-uniform making it limited for quantitative analysis. The extent to which this impacts measurement of sample properties depends strongly on the type of reconstruction. The uniformity in intensities has very little impact on strain measurements. For many ACOM experiments, the intensity of reflections is not weighted strongly when making matches from an existing library of structures. In fact this is the default in py4DSTEM. However, when mapping multiple phases from the same field of view, intensity values can be helpful for separating similar phases and for probing subtle structural changes due to defects (Diebold et al., 2025; Corrêa et al., 2024). It may be possible to incorporate the asymmetry in intensities in ACOM simulations for better matching.
To improve reliability and robustness of MAPED and extend it to a larger tilt series for more accurate quantification, we anticipate the need for automation of the data collection. This requires scripting to automatically tilt the beam, re-center the unscattered beam if necessary, and acquire and save the datasets with sufficient metadata. Acquiring sequential precession datasets increases the memory required to store data, underscoring the importance of detector manufacturers including compression schemes for storing data. Lastly, the MAPED pipeline would be improved through more automated methods for data processing, which will become increasingly important if the number of tilted scans increases.
@@ 5 Conclusion
Analysis and processing of MAPED data are feasible using open-source tools. We expect this method to expand access to precession-enhanced experiments, particularly for challenging samples where conventional techniques may fail. Although we used only five tilts here, our simulations indicate that using more tilts can improve results. In future work, automation of beam tilting and data collection will be essential to scale this sequential precession approach to larger tilt series and make it broadly accessible.

### v2

@@ 3.2 Orientation mapping
Figure 3: (a) dark-field image of polycrystalline aluminum film. (b–e) Conventional diffraction patterns from 4 spots in (a). (f–i) Corresponding MAPED diffraction patterns show enhanced peaks at high scattering angles and reduction of Kikuchi diffraction lines. ACOM maps and dilation ($\epsilon_{xx}$ + $\epsilon_{yy}$) from (j–k) conventional and (l–m) MAPED data with (n) legend
. Dark-field region in (a) marked with rectangle in (j).
To further explore the application of our MAPED method to study material structure, we performed ACOM analysis on a polycrystalline aluminum film (Fig. 3). From the dark-field image (Fig. 3a) we plot four diffraction patterns from the same position. The top row presents conventional STEM diffraction patterns, while the bottom row shows data from the same positions but collected with our MAPED method. As with the data in Fig. 2, incorporating more tilts leads to diffraction information at higher scattering angles. Moreover, the MAPED method suppressed dynamical artifacts such as Kikuchi diffraction, which is seen comparing Fig. 3b&f.
The orientation maps and strain calculations from the ACOM maps are shown in Figs. 3j–m. The conventional diffraction data approach yields high-quality orientation matches. The in-plane rotation maps plus the out-of-plane orientation (Fig. S.9) show most of the grains within the film are approximately aligned with the [111] parallel to the beam with in-plane rotation between grains. Finally, using the structure file as a reference, the strain in each grain is calculated, with the dilation ($\epsilon_{xx}$ + $\epsilon_{yy}$) plotted in Fig. 3k. The full strain maps are shown in Fig. S.11.
The same experiment performed on the identical field of view using MAPED yields qualitatively similar results, but with significantly improved signal-to-noise in both the ACOM maps and strain calculations (Figs. 3l–m, S.10, and S.12). The improvement is especially evident in the strain maps, where the MAPED data produces far more physical results. This trend is evident, for example, in the large grain at the bottom left of the field, highlighted by the arrow in Figs. 3j–m. While the polycrystalline aluminum film studied here serves as a model system to compare ACOM with and without MAPED, we anticipate that this improvement in ACOM matching will be even more significant in challenging samples.
@@ 3.3 Quantification of intensities
Figure 4: Dynamical diffraction simulations to explore impact of few tilt precession on intensity quantification. (a) Si [011] diffraction pattern. The ($\overline{2}00$) and ($\overline{11}1$) intensity values are shown in (b) and (c) as a function of thickness and number of tilts. Intensity values normalized to show percent difference from the intensity value with the highest number of tilts.
Although our MAPED method can increase the signal from high scattering angles as shown in Figs. 2&3, there is an anisotropic nature to the method, which non-uniformly boosts the signal in the diffraction patterns. For example, consider the five diffraction patterns from the silicon sample shown in Fig. 1a. The tilt of the Ewald sphere enhances the signal of some disks more than others.
To better quantify the impact our MAPED method has on diffraction pattern intensities, we performed dynamical Bloch wave simulations in py4DSTEM. In Fig. 4, we plot the the percent difference in intensity in each reflection as compared to the maximum number of tilts (bottom row). As more tilts are included, the differences decrease, and the magnitude of this result will depend on sample thickness, tilt angle, and the number of tilts used. One can imagine how continuously precessing systems, which more finely sample the rotation space, facilitate predictable intensity values and more complete sampling of diffraction space.
It is also important to note how the intensity of these values varies with thickness, which is most easily observed in the intensity plots shown in Fig. S.13. When only one tilt is acquired, as in conventional experiments, the intensity of the ($\overline{11}1$) reflection oscillates with thickness, a well-known feature of dynamical diffraction; this is the first row in the plot in Fig. S.13. With just four additional tilts, as in the MAPED data, these thickness oscillations are significantly decreased and less sensitive to thickness, as with conventional precession experiments (Midgley and Eggeman, 2015). The ($\overline{2}00$) reflection is a kinematically forbidden reflection, and it is expected that with precession that this intensity will be significantly decreased and less sensitive to thickness (Vincent and Midgley, 1994; Ciston et al., 2008; Eggeman et al., 2010). Our simulated MAPED data is in agreement, as including even four additional tilts suppresses the intensity of these reflections. Overall these results highlight how MAPED experiments make intensities more kinematic-like in thicker samples.
@@ 4 Discussion
Precession is well-understood to improve diffraction data by reducing dynamical diffraction artifacts and collecting higher scattering angles. Here, we have highlighted how collecting data at a few tilts can lead to high-quality diffraction data with many of the benefits of integrated precession experiments. The flexibility in choice of cameras and microscopes to collect precession data also poses a significant advantage for 4D-STEM experiments.
Our MAPED method helps with one of the challenges in aligning precession tools for non-integrated cameras. Precession systems that are integrated with a camera address a core challenge of precession 4D-STEM: a precession-induced “descan,” which is a shift of unscattered beam in the back focal plane as a consequence of tilting the incident illumination. This precession-descan is in addition to the common descan observed when an electron beam is rastered across a specimen, and it is elegantly handled by automated routines in the integrated systems. If one does not have a commercial precession integrated with the camera of choice, it is possible to use a camera that is not integrated by having one scan generator do the X-Y scanning in synchronization with the camera and a second scan generator do precession and precession-descan correction. This mode does not have automated routines, so fine tuning of the precession-descan correction with the second scan generator is challenging. Without perfect alignment, the Bragg disks can become blurred due to precession-descan, leaving no ability to correct computationally after acquisition. However, in our MAPED method, we are able to preserve sharp features in our diffraction disks, even in the challenging case of bullseye apertures as shown in Fig. 2, because both types of descan are addressed in post-processing.
@@
An additional consideration is the practicality of acquiring these few tilt 4D-STEM scans. Collecting and processing MAPED datasets requires additional microscope time and computational resources over conventional experiments. However, automated precession pipelines through vendor hardware also require additional overhead and ultimately require similar collection times when competing with fast cameras. An additional practical advantage of the described sequential approach is that the microscope stays in STEM mode, as compared to commercial systems, which usually use a converged probe in TEM mode. Collected data in STEM mode allows easier correlative acquisition of other microscopy signals, including annular dark-field STEM and STEM spectroscopy for multimodal analysis.
Despite these benefits there are limitations for the MAPED precession approach. As studied in Sec. 3.3, especially for thicker samples, the intensities depend on tilts. The extent to which this impacts measurement of sample properties depends strongly on the type of reconstruction. The uniformity in intensities has very little impact on strain measurements. For many ACOM experiments, the intensity of reflections is not weighted strongly when making matches from an existing library of structures. In fact this is the default in py4DSTEM. However, when mapping multiple phases from the same field of view, intensity values can be helpful for separating similar phases and for probing subtle structural changes due to defects (Diebold et al., 2025; Corrêa et al., 2024). It would be possible to incorporate the asymmetry in intensities that comes from discrete instead of continuous tilts in ACOM simulations for better matching. The practical benefits and limitations of the MAPED approach, as compared to conventional precession, could be more directly compared on a microscope with built-in precession.
To improve reliability and robustness of MAPED and extend it to a larger tilt series for more accurate quantification, we anticipate the need for automation of the data collection. This requires scripting to automatically tilt the beam, re-center the unscattered beam if necessary, and acquire and save the datasets with sufficient metadata. Acquiring sequential precession datasets increases the memory required to store data, underscoring the importance of compression schemes for storing data. Lastly, the MAPED pipeline would be improved through more automated methods for data processing, which will become increasingly important if the number of tilted scans increases.
@@ 5 Conclusion
Analysis and processing of MAPED data are feasible using open-source tools. We expect this method to expand access to precession-enhanced experiments, particularly for challenging samples where conventional techniques may fail. Although we used only five tilts here, our simulations indicate that using more tilts can improve results. In future work, automation of beam tilting and data collection will be essential to scale this sequential precession approach to larger tilt series and make it more broadly accessible.